# Randomization Tests for Conditional Group Symmetry

### v1

@@ Abstract
Symmetry plays a central role in the sciences, machine learning, and statistics. While statistical tests for the presence of distributional invariance with respect to groups have a long history, tests for conditional symmetry in the form of equivariance or conditional invariance are absent from the literature. This work initiates the study of nonparametric randomization tests for symmetry (invariance or equivariance) of a conditional distribution under the action of a specified locally compact group. We develop a general framework for randomization tests with finite-sample Type I error control and, using kernel methods, implement tests with finite-sample power lower bounds. We also describe and implement approximate versions of the tests, which are asymptotically consistent. We study their properties empirically on synthetic examples, and on applications to testing for symmetry in two problems from high-energy particle physics.
62G10,
62H15,
62H05,
62P35,
equivariance,
invariance,
symmetry,
hypothesis test,
randomization test,
@@ 1 Introduction
Symmetry plays a central role in science and engineering, and appears in numerous statistical problems both classical (Lehmann and Casella, 1998; Lehmann and Romano, 2005) and modern (e.g., Cohen, Geiger and Weiler, 2019; Chen, Dobriban and Lee, 2020; Huang, Orbanz and Austern, 2022; Bronstein et al., 2021). There is an extensive literature on hypothesis tests for symmetry in the form of invariance under a group of transformations, dating back at least to the work of Hoeffding (1952), which generalized older ideas based on testing permutation-invariance (Fisher, 1935; Pitman, 1937). More recent developments in the area pertain to randomized versions of tests for invariance (Romano, 1988, 1989; Hemerik and Goeman, 2017; Hemerik and Goeman, 2018; Dobriban, 2022; Ramdas et al., 2023; Koning and Hemerik, 2024) or approximate invariance (Canay, Romano and Shaikh, 2017); see Ritzwoller, Romano and Shaikh (2024) for a recent review. One takeaway from that line of work is that group-based randomization tests have finite-sample Type I error control (Hoeffding, 1952), power against alternatives (Romano, 1989; Dobriban, 2022), and achieve various notions of optimality and robustness (Romano, 1990; Kashlak, 2022; Koning and Hemerik, 2024).
@@
The literature on group-based randomization tests deals with the invariance of an unknown probability distribution $P_{X}$ from which samples $X_{1},\dotsc,X_{n}$ are obtained. In this paper, we initiate the study of hypothesis tests for the symmetry of a conditional distribution, $P_{Y|X}$, to be tested from samples $(X_{1},Y_{1}),\dotsc,$ $(X_{n},Y_{n})$. Under certain conditions, this can be reduced to a problem for which group-invariance tests are applicable, but those conditions are not satisfied in many settings. The primary motivation for conditional tests is the central role played by symmetry in modern science (Rosen, 1995; Gross, 1996) and the increasingly data-driven nature of scientific discovery, particularly in settings in which the relationships between multiple objects are a key feature (Kofler and Schlötterer, 2012; ATLAS Collaboration, 2017; Collins, Howe and Nachman, 2018; Cristadoro, Degli Esposti and Altmann, 2018; Chmiela et al., 2018; Karagiorgi et al., 2022; Birman et al., 2022; Yim et al., 2023; Watson et al., 2023). A secondary motivation is the prominence of symmetry in certain prediction problems from machine learning (Bronstein et al., 2021).
@@
Figure 1: First column: a random sample from a marginal distribution on $\mathbb{R}^{2}$. A marginal distribution is invariant with respect to rotations about the origin if rotating the distribution does not change it. Second and third columns: conditional distributions $P_{Y|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)$ for $Y\in\mathbb{R}^{2}$ (triangle) given three different values of $x\in\mathbb{R}^{2}$ (circle), matched by color. A conditional distribution is rotationally-equivariant if rotations of $x$ transform the distribution of $Y|X=x$ with a corresponding rotation, and conditionally invariant if rotations of $x$ do not change the distribution of $Y|X=x$.
@@
Figure 1 illustrates the differences between marginal and conditional symmetries with a simple example. Whereas symmetry of a probability distribution manifests only as invariance under a group of transformations, a conditional distribution may exhibit a form of symmetry that only appears in relationships between variables, known as equivariance: a transformation of $X$ induces a corresponding transformation of $Y$. As a consequence, testing for conditional symmetry presents certain challenges absent in tests for (marginal) distributional symmetry. In the latter, all tests can be understood as variations on averaging over the group with respect to the “uniform distribution”, a property that stems from the fact that every invariant distribution can be factored into the product of the group’s uniform distribution and a distribution corresponding to an independent random variable. Conditional symmetries do not have such a neat and convenient characterization, in part because they may hold with respect to a non-compact group, for which a uniform distribution does not exist. This is in contrast to distributional invariance, which may hold only with respect to a compact group. Beyond the technical challenges of non-compact groups, there are additional statistical and computational challenges to overcome.
One contribution of the present work is a number of characterizations of conditional symmetry that can be used as the basis for a hypothesis test. The first, which generalizes in practically important ways a result of Bloem-Reddy and Teh (2020), characterizes conditional symmetry as a certain conditional independence relationship. However, generic tests for conditional independence that control Type I error over the entire class of conditionally independent distributions are known to have poor statistical properties (Shah and Peters, 2020). To that end, we establish further characterizations of conditional symmetry that can be used as the basis for group-based conditional randomization tests (Candes et al., 2018), which restrict the null hypothesis class by conditioning on part of the data and therefore have much better statistical properties than the aforementioned conditional independence tests.
Related work.
To our knowledge, the existing literature does not contain any general-purpose tests for conditional symmetry. In addition to typical hypothesis testing applications, such a tool would also be of interest as model-checking criteria in prediction applications where, for example, there is reason to believe that transforming the input in a particular way is equivalent to a corresponding transformation in the output. Prediction problems with symmetry have been a subject of intense study in machine learning; Bronstein et al. (2021) provide a recent review. Besides the literature on group-based randomization tests, some recent methods in machine learning aim to directly learn or estimate symmetry groups from data (Krippendorf and Syvaeri, 2020; Zhou, Knowles and Finn, 2021; Desai, Nachman and Thaler, 2022; Dehmamy et al., 2021; Yang et al., 2023). They do not address the corresponding inferential questions, and the present work can be viewed as a complement to those estimation methods.
To the best of our knowledge, the only work that addresses any notion of testing for conditional symmetry is that of Christie and Aston (2023), who proposed tests for the invariance of the conditional expectation $f(x)=\mathbb{E}[Y|X=x]$, $f\colon\mathbf{X}\to\mathbb{R}$, under the action of some compact group. [^1] We note that the primary aim of Christie and Aston (2023) is to estimate the largest group under which $f$ is invariant, which amounts to performing a collection of tests over a subgroup lattice of some candidate largest group. In principle, our tests could be substituted into their procedure, though we do not address that problem here. Those authors’ focus on testing for the invariance of a function leads them in a different direction, and their tests are limited in their applicability compared to the tests proposed here.
Contributions.
@@
The present work formulates a general framework for testing the invariance or equivariance of a conditional distribution under the action of a specified locally compact group $\mathbf{G}$, based on a single independent and identically distributed (i.i.d.) sample. These tests can be used to detect the presence or absence of a particular conditional symmetry in data, with that symmetry specified by hypothesis or by a data-driven method that has learned the symmetry. In situations with known or assumed symmetry, these tests may also be used as model-checking criteria for models meant to exhibit that symmetry.
@@
In addition to a general formulation of symmetry-based conditional randomization tests for which we prove Type I error control properties, we provide specific kernel-based instantiations of our tests, which allow these tests to be used on any data type for which a kernel can be defined. Using a test statistic based on kernel maximum mean discrepancy, we derive a finite-sample lower bound on the test’s power against non-symmetric alternatives, which also implies that the test is consistent as the number of observations $n\to\infty$.
We demonstrate the application of our tests on synthetic data, and on two problems from high-energy particle physics. We also demonstrate the tests’ use for measuring the symmetry properties of a MNIST classifier that is trained to be invariant by using data augmentation. Computer code required to run the experiments can be found on an associated GitHub repository. [^2] https://github.com/chiukenny/Tests-for-Conditional-Symmetry
Outline.
@@
The paper is structured as follows. In Section 2, we review the necessary basics of groups (additional technical aspects are contained in Section A.1), and give an overview of randomization tests for group invariance. We also fill a gap in the literature on such tests, showing that they are conditional level-$\alpha$ tests for any compact group that acts continuously (Theorem 2.2). In Section 3, we introduce the framework for conditional symmetry and describe some examples. Section 4 is the main technical section, in which conditional symmetry is characterized via conditional independence and certain distributional identities. We propose a general framework for conditional randomization tests of symmetry based on those results in Section 5. In Section 6, we describe specific nonparametric tests based on kernel maximum mean discrepancy, along with finite-sample lower bounds on power against non-symmetric alternatives. Section 7 contains experiments on synthetic data and on particle physics data, as well as a demonstration of how the tests can be used to assess goodness-of-fit for flexible prediction models meant to exhibit symmetry.
@@ 2 Background
Before reviewing the basics of groups and randomization tests for group invariance, we briefly introduce some technical necessities and notation. Throughout, we use $\mathbf{X}$ to denote a topological space and $\mathbf{S}_{\mathbf{X}}$ its Borel $\sigma$-algebra, so that $(\mathbf{X},\mathbf{S}_{\mathbf{X}})$ is a standard Borel measurable space. Unless stated otherwise, we refer to $\mathbf{X}$ as a measurable space, and likewise for $\mathbf{Y}$ and $\mathbf{M}$. Let $\mathcal{P}(\mathbf{X})$ denote the set of all probability measures on $\mathbf{X}$. For $P\in\mathcal{P}(\mathbf{X})$, we write $X\sim P$ for a random variable $X$ sampled from $P$. A vector or tuple of random variables is denoted by $X_{1:n}\coloneq(X_{1},\dotsc,X_{n})$. For any measurable function $f$ on $\mathbf{X}$, let $f_{*}P$ denote the pushforward measure with $f_{*}P(A)=P(f^{-1}(A))$ for all measurable sets $A\in\mathbf{S}_{\mathbf{X}}$. We use $\delta_{x}$ to denote the Dirac measure at a point $x$, and =$\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}$ to denote equality in distribution. Let $[n]$ denote the set $\{1,2,\dotsc,n\}$.
@@ 2.1 Groups and group actions
This work focuses on equivariant symmetries that can be encoded in the language of group theory, which comprise many of the symmetries of interest in modern statistical, computational, and scientific problems. We assume a basic familiarity with groups and group actions; see Appendix A for background on groups. A thorough treatment of group theory and its applications in statistical contexts can be found in (Eaton, 1989; Wijsman, 1990).
@@
We denote a group by $\mathbf{G}$ and its elements by $g$. For convenience, group composition is written multiplicatively, i.e, $g_{1}g_{2}=g_{1}\cdot g_{2}$. We assume that $\mathbf{G}$ has a topology that is locally compact, second countable, and Hausdorff (lcscH), and which makes the group operations continuous. We may then take $\mathbf{S}_{\mathbf{G}}$ as the Borel $\sigma$-algebra, making $\mathbf{G}$ a standard Borel space. For $A\subseteq\mathbf{G}$ and $g\in\mathbf{G}$, we write $gA=\{gh:h\in A\}$ and $Ag=\{hg:h\in A\}$. A measure $\nu$ on $\mathbf{G}$ is said to be left-invariant if $\nu(gA)=\nu(A)$ for all $A\in\mathbf{S}_{\mathbf{G}}$, and right-invariant if $\nu(Ag)=\nu(A)$. When $\mathbf{G}$ is lcscH, there exist left- and right-invariant $\sigma$-finite measures $\lambda_{\mathbf{G}}$ and $\tilde{\lambda}_{\mathbf{G}}$, known as left- and right-Haar measures, respectively, that are unique up to scaling (Folland, 2016, Ch. 2.2). We use $\lambda$ to denote left-Haar measure. If $\mathbf{G}$ is compact, then $\lambda=\tilde{\lambda}$, and the unique normalized Haar measure acts as the uniform probability measure over the group.
A group $\mathbf{G}$ acts measurably on a set $\mathbf{X}$ if the group action $(g,x)\mapsto gx$ is a measurable function. Throughout, all actions are assumed to be continuous (and therefore measurable), and for convenience we simply say that $\mathbf{G}$ acts on $\mathbf{X}$. For a set $A\subseteq\mathbf{X}$, $gA=\{gx:x\in A\}$. The stabilizer subgroup of $x\in\mathbf{X}$ is the subset of group actions that leave $x$ invariant, denoted $\mathbf{G}_{x}=\{g\in\mathbf{G}:gx=x\}$. The action is free (or exact) if $gx=x$ implies that $g=\textrm{id}$, in which case $\mathbf{G}_{x}=\{\textrm{id}\}$ for all $x\in\mathbf{X}$. The $\mathbf{G}$-orbit of $x\in\mathbf{X}$ is the subset of all points in $\mathbf{X}$ reachable through a group action in $\mathbf{G}$, denoted $O(x)=\{gx:g\in\mathbf{G}\}$. Each orbit is an equivalence class of $\mathbf{X}$, where two points are equivalent if and only if they belong to the same orbit. If $\mathbf{X}$ has only one orbit, then the action is transitive.

### v2

@@ Abstract
Symmetry plays a central role in the sciences, machine learning, and statistics. While statistical tests for the presence of distributional invariance with respect to groups have a long history, tests for conditional symmetry in the form of equivariance or conditional invariance are absent from the literature. This work initiates the study of nonparametric randomization tests for symmetry (invariance or equivariance) of a conditional distribution under the action of a specified locally compact group. We develop a general framework for randomization tests with finite-sample Type I error control and, using kernel methods, implement tests with finite-sample power lower bounds. We also describe and implement approximate versions of the tests, which are asymptotically consistent. We study their properties empirically using synthetic examples and applications to testing for symmetry in two problems from high-energy particle physics.
@@ 1 Introduction
Symmetry plays a central role in science and engineering, and appears in numerous statistical problems, both classical (Lehmann and Casella, 1998; Lehmann and Romano, 2005) and modern (e.g., Cohen, Geiger and Weiler, 2019; Chen, Dobriban and Lee, 2020; Huang, Orbanz and Austern, 2022; Bronstein et al., 2021). There is an extensive literature on hypothesis tests for symmetry in the form of distributional invariance under a group of transformations, dating back at least to the work of Hoeffding (1952), which generalized older ideas based on testing permutation-invariance (Fisher, 1935; Pitman, 1937). More recent developments in the area pertain to randomized versions of tests for invariance (Romano, 1988, 1989; Hemerik and Goeman, 2017; Hemerik and Goeman, 2018; Dobriban, 2022; Ramdas et al., 2023a; Koning and Hemerik, 2024) or approximate invariance (Canay, Romano and Shaikh, 2017); see Ritzwoller, Romano and Shaikh (2024) for a recent review. One takeaway from that line of work is that group-based randomization tests have finite-sample Type I error control (Hoeffding, 1952), power against alternatives (Romano, 1989; Dobriban, 2022), and achieve various notions of optimality and robustness under appropriate conditions (Romano, 1990; Kashlak, 2022; Kim, Balakrishnan and Wasserman, 2022; Koning and Hemerik, 2024).
@@
The literature on group-based randomization tests deals with the invariance of an unknown probability distribution $P_{X}$ from which samples $X_{1},\dotsc,X_{n}$ are obtained. In this paper, we initiate the study of hypothesis tests for the symmetry of a conditional distribution, $P_{Y|X}$, to be tested from samples $(X_{1},Y_{1}),\dotsc,$ $(X_{n},Y_{n})$. The primary motivation for conditional tests is the central role played by symmetry in modern science (Rosen, 1995; Gross, 1996) and the increasingly data-driven nature of scientific discovery, particularly in settings in which the relationships between multiple objects are a key feature (Kofler and Schlötterer, 2012; ATLAS Collaboration, 2017; Collins, Howe and Nachman, 2018; Cristadoro, Degli Esposti and Altmann, 2018; Chmiela et al., 2018; Karagiorgi et al., 2022; Birman et al., 2022; Yim et al., 2023; Watson et al., 2023). A secondary motivation is the prominence of symmetry in certain prediction problems from machine learning; Bronstein et al. (2021) provide a recent review.
@@
Figure 1: Column 1: a random sample from a distribution on $\mathbb{R}^{2}$. A distribution is invariant with respect to rotations about the origin if rotating the distribution does not change it. Columns 2 and 3: conditional distributions $P_{Y|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)$ for $Y\in\mathbb{R}^{2}$ (triangle) given three different values of $x\in\mathbb{R}^{2}$ (circle), matched by color. A conditional distribution is rotationally-equivariant if rotations of $x$ transform the distribution of $Y|X=x$ with a corresponding rotation, and conditionally invariant if rotations of $x$ do not change the distribution of $Y|X=x$.
@@
Figure 1 illustrates the differences between marginal and conditional symmetries with a simple example. Whereas symmetry of a probability distribution manifests as invariance under a group of transformations, a conditional distribution may exhibit a form of symmetry that only appears in relationships between variables, known as equivariance: a transformation of $X$ induces a corresponding transformation of $Y$. Under certain conditions, detecting equivariance can be reduced to a problem for which existing group-invariance tests are applicable, but those conditions are not satisfied in many settings. Tests for (marginal) distributional group symmetry are based on the fact that an invariant distribution factors into the independent product of the group’s “uniform distribution” and a distribution over group-induced equivalence classes known as orbits—a characterization that follows from the fact that marginal invariance can only hold with respect to compact groups. The independence between group transformation and orbit is key to the computational and statistical simplicity of tests for marginal invariance. In contrast, conditional symmetries do not have such a convenient characterization, in part because they may hold with respect to a non-compact group, for which a uniform distribution does not exist. As a consequence, designing tests for conditional symmetry requires overcoming new technical and statistical challenges.
Contributions
@@
The present work initiates the study of hypothesis testing for conditional group symmetry and formulates a general framework for such tests based on a single independent and identically distributed (i.i.d.) sample. In order to do so, we develop a number of novel characterizations of the invariance and equivariance of a conditional distribution under the action of a locally compact group. The tests we develop can be used to detect the presence or absence of a particular conditional symmetry in data, with that symmetry specified by hypothesis. The tests also may be used as model-checking criteria for models meant to exhibit a symmetry.
@@
In addition to a general formulation of symmetry-based conditional randomization tests for which we prove Type I error control properties, we provide specific kernel-based instantiations of our tests, which allow the tests to be used on any data type for which a kernel can be defined. The testing procedure approximates the conditional distribution of the estimated maximum mean discrepancy between two samples from a suitable null hypothesis distribution, conditioned on certain symmetry-induced statistics of the observations. Using this test, we derive a finite-sample lower bound on the test’s power against non-symmetric alternatives, which also implies that the test is asymptotically consistent as the number of observations $n\to\infty$. We explore the properties of our tests empirically on synthetic data and demonstrate their application to two problems from high-energy particle physics.
Related work
To our knowledge, general-purpose tests for conditional symmetry have not appeared in the existing literature. The focus of this paper is on tests for conditional group symmetry, which should not be confused with the literature that tests conditional distributions for symmetry about zero (or some other value) for every value of the conditioning variable (e.g., Zheng, 1998; Bai and Ng, 2001; Su, 2006). In contrast to the present work, those tests do not consider symmetry in the relationship between the variables and are applicable to a disjoint set of problems from those considered here. All mentions of conditional symmetry in this work will refer to the group invariance or equivariance of a conditional distribution, as we define in Section 2.3.
Some recent methods in machine learning aim to estimate symmetry groups from data (Krippendorf and Syvaeri, 2020; Zhou, Knowles and Finn, 2021; Desai, Nachman and Thaler, 2022; Dehmamy et al., 2021; Yang et al., 2023). They do not address the corresponding inferential questions, and the present work can be viewed as a complement to those estimation methods.
The only work that we are aware of that addresses any notion of testing for conditional symmetry is that of Christie and Aston (2025), who proposed tests for the invariance of the conditional expectation $f(x)=\mathbb{E}[Y|X=x]$, $f\colon\mathbf{X}\to\mathbb{R}$, under the action of a specified group. [^1] We note that the primary aim of Christie and Aston (2025) is to estimate the largest group under which $f$ is invariant, which amounts to performing a collection of tests over a subgroup lattice of some candidate largest group. In principle, our tests could be substituted into their procedure, though we do not address that problem here. Testing for the invariance of the conditional expectation is not the same as testing the invariance of the conditional distribution, and those authors’ focus on the former objective leads them to formulate tests that rely on certain regularity properties of $f$ (e.g., smoothness), for which there are no distributional analogues. In contrast, the tests that we develop here are applicable to a different set of problems: testing for distributional symmetry, and with theoretical properties that do not rely on regularity of the conditional moments. Moreover, the statistical properties of the tests in (Christie and Aston, 2025) are sensitive to the choice of a number of user-specified parameters. In particular, a sampling distribution on the group is chosen by heuristic. In the tests we develop, the required sampling distribution over the group and other test parameters are estimated from data in such a way that preserves the validity of the test while improving power against alternatives.
The methods we develop rely on group-theoretic inversion probability kernels, which have not been used widely in statistics or machine learning, likely owing to their relatively recent appearance in probability (Kallenberg, 2011). Since their appearance in a preliminary version of this work (Chiu and Bloem-Reddy, 2023), inversion kernels have been used to solve problems involving symmetry in machine learning (Cornish, 2024; Lawrence et al., 2025; Zhang et al., 2025). Using inversion kernels, we characterize conditional symmetry through a particular conditional independence relation, generalizing a result of Bloem-Reddy and Teh (2020) that holds in a much more limited setting. Those authors used the result to study equivariant neural networks and did not consider testing for conditional symmetry. Deterministic inversions have been used more widely in the machine learning literature (e.g., Bloem-Reddy and Teh, 2020; Winter et al., 2022; Kaba et al., 2023), and have counterparts in classical statistical problems involving groups (Fraser, 1961; Eaton, 1989; Eaton and Sudderth, 1999).
Outline
@@
In Section 2, we review the necessary basics of groups (additional technical aspects are contained in Section A.1), give an overview of randomization tests for group invariance, and introduce the framework for conditional symmetry along with some examples. Section 3 is the main technical section, in which conditional symmetry is characterized via conditional independence and certain distributional identities. We propose a general framework for conditional randomization tests of symmetry based on those results in Section 4. In Section 5, we describe specific nonparametric tests based on kernel maximum mean discrepancy, along with finite-sample lower bounds on power against non-symmetric alternatives. Section 6 contains experiments that apply our tests to synthetic data and to particle physics data.
@@ 2 Background
We first introduce some technical necessities. Throughout, we use $\mathbf{X}$ to denote a topological space and $\mathbf{S}_{\mathbf{X}}$ its Borel $\sigma$-algebra, so that $(\mathbf{X},\mathbf{S}_{\mathbf{X}})$ is a standard Borel measurable space. Unless stated otherwise, we refer to $\mathbf{X}$ as a measurable space, and likewise for $\mathbf{Y}$, $\mathbf{M}$, and $\mathbf{R}$ (distinct from $\mathbb{R}$, the space of real numbers). Let $\mathcal{P}(\mathbf{X})$ denote the set of all probability measures on $\mathbf{X}$. For $P\in\mathcal{P}(\mathbf{X})$, we write $X\sim P$ for a random variable $X$ sampled from $P$. A vector or tuple of random variables is denoted by $X_{1:n}\coloneq(X_{1},\dotsc,X_{n})$. For any measurable function $f$ on $\mathbf{X}$, let $f_{*}P$ denote the pushforward measure with $f_{*}P(A)=P(f^{-1}(A))$ for all measurable sets $A\in\mathbf{S}_{\mathbf{X}}$. We use $\delta_{x}$ to denote the Dirac measure at a point $x$, and $\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}$ to denote equality in distribution. Let $[n]$ denote the set $\{1,2,\dotsc,n\}$.
@@ 2.1 Groups and group actions
This work focuses on symmetries that can be encoded in the mathematics of group theory, which comprise many of the symmetries of interest in modern statistical, computational, and scientific problems. We assume a basic familiarity with groups and group actions; see Appendix A for background. Thorough treatments of classical group theoretic applications in statistics, along with relevant mathematical details, can be found in (Eaton, 1989; Wijsman, 1990).
@@
We denote a group by $\mathbf{G}$ and its elements by $g$. For convenience, group composition is written multiplicatively, i.e, $g_{1}g_{2}=g_{1}\cdot g_{2}$. We assume that $\mathbf{G}$ has a topology that is locally compact, second countable, and Hausdorff (lcscH), and which makes the group operations continuous. We may then take $\mathbf{S}_{\mathbf{G}}$ as the Borel $\sigma$-algebra, making $\mathbf{G}$ a standard Borel space. When $\mathbf{G}$ is lcscH, there exist left- and right-invariant $\sigma$-finite measures $\lambda_{\mathbf{G}}$ and $\tilde{\lambda}_{\mathbf{G}}$, known as left- and right-Haar measures, respectively, that are unique up to scaling (Folland, 2016, Ch. 2.2). For convenience, we omit the subscript and use $\lambda$ to denote left-Haar measure. If $\mathbf{G}$ is compact, then $\lambda=\tilde{\lambda}$, and the unique normalized Haar measure acts as the uniform probability measure over the group.
A group $\mathbf{G}$ acts measurably on a set $\mathbf{X}$ if the group action $(g,x)\mapsto gx$ is a measurable function. Throughout, all actions are assumed to be continuous (and therefore measurable), which we imply by saying that $\mathbf{G}$ acts on $\mathbf{X}$. For a set $A\subseteq\mathbf{X}$, $gA=\{gx:x\in A\}$. The $\mathbf{G}$-orbit of $x\in\mathbf{X}$ is the subset of all points reachable through an action in $\mathbf{G}$, denoted $O(x)=\{gx:g\in\mathbf{G}\}$. If $\mathbf{X}$ has only one orbit, then the action is transitive. On each orbit $O(x)$, we can choose an element $r_{x}\in O(x)$ as the orbit representative, so that $r_{x}=gx$ for some $g\in\mathbf{G}$. The element $r_{x}$ can be thought of as an “origin” on the orbit. For a particular choice of representatives, the subset of $\mathbf{X}$ consisting of each orbit’s representative is denoted by $\mathbf{R}$. A generic element of $\mathbf{R}$ is denoted by $r$, whereas $r_{x}$ indicates the element of $\mathbf{R}$ on the orbit of $x$, i.e., $\mathbf{R}\cap O(x)=\{r_{x}\}$. A function $\rho\colon\mathbf{X}\to\mathbf{R}$ that maps elements of $\mathbf{X}$ onto their corresponding orbit representatives is called an orbit selector; that is, $\rho(x)=r_{x}$. Any orbit selector makes for a useful choice of a maximal invariant, which more generally is a $\mathbf{G}$-invariant function $M\colon\mathbf{X}\to\mathbf{M}$ that takes a different value on each orbit, so that $M(x)=M(x^{\prime})$ if and only if $gx=x^{\prime}$ for some $g\in\mathbf{G}$. Maximal invariants are not unique; we use $M$ to denote a generic maximal invariant. Figure 2 depicts these quantities for $\mathrm{SO}(d)$—the compact group of $d$-dimensional rotations about the origin—acting on $\mathbb{R}^{d}$, for $d=2$ and $d=3$. (Figure 2 also depicts representative inversions $\psi^{\circ}$}, which are defined in Section 3.) The stabilizer subgroup of $x\in\mathbf{X}$ is the subset of group actions that leave $x$ invariant, denoted $\mathbf{G}_{x}=\{g\in\mathbf{G}:gx=x\}$. The group action is free if $\mathbf{G}_{x}=\{\textrm{id}\}$ for all $x\in\mathbf{X}$.
@@
[…]
Figure 2: Left: a graphical depiction of a point $x\in\mathbb{R}^{2}$, its orbit $O(x)=\{x^{\prime}\in\mathbb{R}^{2}:\|x^{\prime}\|_{2}=M(x)\}$, its orbit representative $r_{x}\in\mathbf{R}=\{(z,0):z\geq 0\}$, and the role of the orbit selector $\rho$, maximal invariant $M$, and representative inversion $\psi^{\circ}$} under the action of $\mathrm{SO}(2)$. Right: a graphical depiction of a point $x\in\mathbb{R}^{3}$, its orbit $O(x)$ and orbit representative $r_{x}\in\mathbf{R}$ under the action of $\mathrm{SO}(3)$. Depicted are two rotations $g_{1}$ and $g_{2}$ that take $r_{x}$ to $x$. The inversion kernel $P_{\Psi|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)$ is the uniform distribution over $\{g\in\mathrm{SO}(3):x=gr_{x}\}$.